A competitive firm has the short-run cost function C(y) = 12y3  - 8y2+30y+12. At what price will the firm agree to produce in the short-run? What is the shutdown condition for this firm?

 



1
Expert's answer
2021-12-01T21:58:29-0500

Let the selling price be p

"TR=p\times y\\TR=py"

We can derive MR from this, "MR=\frac{dTR}{dy}=p"

"TC=C(y) = 12y\\^3−8y\\^2+30y+12"

Now, "MC=\frac{dTC}{dy}=48y\\^2-16y+30"

The firm will agree to produce in the short run when "MR>MC p>48y2-16y+30"  The shutdown condition for this firm is when "MR<MCp<48y2-16y+30"